# Hilbert properties of varieties

### v1

@@ Acknowledgements
The authors are grateful to Finn Bartsch, Lior Bary-Soroker, Pietro Corvaja, Pierre Dèbes, Julian Demeio, Davide Lombardo, Dan Loughran, Cedric Luger, Sebastian Petersen, Sam Streeter, and Umberto Zannier for the exchange of many ideas relevant to this survey. They are indebted to Jean-Louis Colliot-Thélène, Sebastian Petersen, Sam Streeter and Umberto Zannier for numerous helpful remarks and corrections on a preliminary version.
@@ 1.1. Hilbert’s theorem
In geometric terms, this argument of Hilbert exploits that the quotient variety $\mathbb{A}^{n}_{\mathbb{Q}}/S_{n}$ is again a rational variety (i.e., birationally equivalent to $\mathbb{A}^{n}_{\mathbb{Q}}$), in fact $\mathbb{A}^{n}_{\mathbb{Q}}/S_{n}\cong\mathbb{A}^{n}_{\mathbb{Q}}$, so that the Galois cover $\mathbb{A}^{n}_{\mathbb{Q}}\rightarrow\mathbb{A}^{n}_{\mathbb{Q}}/S_{n}$ can be specialized to a Galois extension of $\mathbb{Q}$ with the same Galois group. (We will return to this geometric interpretation in Section 1.3.) In what became known as the Noether program, Emmy Noether [Noe18] started the investigation of whether $\mathbb{A}^{n}_{\mathbb{Q}}/G$ is rational for all finite groups $G\leq S_{n}$, which if true would give a strategy to solve the inverse Galois problem: Every finite group $G$ would be the Galois group of a Galois extension of $\mathbb{Q}$. However, counterexamples found by Swan [Swa69] and Lenstra [Len74] show that $\mathbb{A}^{n}_{\mathbb{Q}}/G$ is non-rational already for the cyclic groups $G=\mathbb{Z}/47\mathbb{Z}$ respectively $G=\mathbb{Z}/8\mathbb{Z}$. (See the references in [Len74] for more on the history of this problem.)
@@ 1.2. Hilbertian fields
Extensive sources on Hilbertian fields are [Lan83, Chapter 9], [FJ08, Chapters 12-16], [MM18, Chapter IV], and [Völ96, Chapter 1]. For $K$ of positive characteristic, there are also variants of Definition 1.8 in the literature that are not equivalent to it. Our definition (more precisely Proposition 1.11(2)) is the one used in [FJ08] from the second edition onwards and is the one commonly used nowadays. There are also many actual strengthenings and weakenings in the literature, and we mention only some of them: fully Hilbertian fields [BP13], $g$-Hilbertian fields [FJ08, Chapter 13.6], RG-Hilbertian fields and R-Hilbertian fields [DH99], real Hilbertian fields [FHV94], and $\Theta$-Hilbertian fields [FH20]. There are also the notions of Hilbertian rings [FJ08, Chapter 13.4] and integrally Hilbertian rings [BD25].
@@ 1.3. The Hilbert property
Motivated by the failure of the Noether program, Colliot-Thélène, Sansuc and Serre [CS87, Ser08] investigated whether it might be possible to specialize the cover $\mathbb{A}_{\mathbb{Q}}^{n}\rightarrow\mathbb{A}_{\mathbb{Q}}^{n}/G$ to a Galois extension of $\mathbb{Q}$ with Galois group $G$ even if $\mathbb{A}^{n}_{\mathbb{Q}}/G$ is not rational. This led them to the following definition. Here, $X$ is a $K$-variety (i.e. an integral separated scheme of finite type over $K$), and a cover is a finite surjective generically étale morphism of $K$-varieties.
@@
The Hilbert property and thin sets are discussed in the textbooks [Ser08, Chapter 3], [Ser97, Chapter 9], and [FJ08, Chapter 13.5]. The first definition of HP in [CS87, p. 189] (there called “Hilbert type”) was phrased in terms of specializations of polynomials, [Ser08] introduces the notion of thin sets, while the definition we give here is adjusted for positive characteristic (cf. Remark 1.13). Some results on HP predate its formulation as a property, like [Man86, Theorem 46.1], whose first Russian edition appeared in 1972.
@@ 1.4. The weak Hilbert property
Question 1.34 ([CZ17, Question-Conjecture 2]).
@@
In fact, [CZ17] ask the question with $X$ only required to be normal, but then the answer is negative, as Remark 1.32 shows. As evidence that the answer to Question 1.34 might be positive, Corvaja and Zannier mention the case of curves (where it holds, see Example 1.30 and Section 3), conjectures of Vojta, and Manin’s conjecture in the case of Fano varieties. Like with Question 1.25, they suggest that the answer might be positive only after allowing a finite extension of the base field. We discuss this in more detail in Section 1.6.
@@ 1.5. Integral Hilbert properties
Let $K$ be a field of characteristic zero and $X$ a normal and geometrically integral $K$-variety. A positive answer to Question 1.34 would mean that if $K$ is a number field, $X$ is smooth and projective, and $X(K)$ is Zariski-dense in $X$, then $X$ has WHP, but Corvaja and Zannier considered the possibility (cf. Remark 1.35) that this might hold only potentially, in the following sense:
@@
So in the case where $K$ is a number field, a positive answer to this weaker version of Question 1.34 would offer a precise prediction when $X$ should have potential WHP, namely if and only if $X(L)$ is Zariski-dense in $X$ for some finite extension $L/K$. We now discuss for which $X$ the latter condition is expected to be satisfied. Results and conjectures in this subsection will only be used to motivate some questions and results later on.
@@
In [Cam04, Definition 2.1] (see also [Cam11, Definition 8.1] or [BJL24, Definition 1.2]) Campana introduced the class of special varieties and he conjectured that the arithmetic properties of special varieties are opposite to those conjectured by Lang for varieties of general type; see [Cam11, Conjecture 13.23]. Combined with the expected positive answer to the weakened version (see Remark 1.35) of Question 1.34, extended to finitely generated fields, this gives:
@@ 2.4. Product and fibration theorems
(1) was first published in [BFP14]; an unpublished manuscript of van den Dries from the 1990’s proves this using nonstandard methods; yet another proof was given in [BFP25, Corollary 3.5]; for $K$ a number field, (1) also follows from [HW16, Lemma 8.12]. (2) follows from [Lug24a, Theorem 1.4], generalizing the special case of smooth proper varieties in [CD+22, Theorem 1.9]. ∎
@@ 3. Curves
Let $K$ be Hilbertian and $X$ a curve over $K$. To see that the statement of Theorem 3.2 does not hold in this generality, recall that if $K$ is also PAC, all such $X$ have HP (Example 1.19), so the implication $\Rightarrow$ in (1)-(4) does not hold. The following Example 3.5 shows that also the implication $\Leftarrow$ in (2) does not hold. However, Proposition 3.1 implies that $\Leftarrow$ in (1) and (3) still hold, and a positive answer to Question 1.46 would imply that also $\Leftarrow$ in (4) holds in general.
@@ 4. Algebraic groups
Since connected algebraic groups are special (Example 1.44), Theorem 4.2(3) confirms a case of Conjecture 1.43. A positive answer to Question 1.46 would imply that (3) holds for arbitrary Hilbertian $K$, which in the case where $G$ is an abelian variety was confirmed for function fields $K$ in [Jav25, Theorem A]. However, Example 3.5 shows that already in the case of elliptic curves, (2) does not hold even for function fields $K$.
@@ 3. $(c)$
(1) is [Liu22, Theorem 1.3], and (2) is [CD+22, Theorem 1.4]. For cyclic $\Gamma$ and $K$ a number field, (1) in the case $G=(\mathbb{G}_{m})^{r}$ and (2) in the case $G=E^{r}$ for an elliptic curve $E$ was proven in [Zan10], which for $G=(\mathbb{G}_{m})^{r}$ built on results from [DZ07] over $\mathbb{Q}^{\rm ab}$.
@@ 5.2. Surfaces with $\kappa(X)=0$
Let $K$ be a number field, let $C$ be a smooth plane cubic in $\mathbb{P}^{2}_{K}$ and let $X=\mathbb{P}_{K}^{2}\setminus C$. Using conic fibrations on $X$, [Coc19, Theorem 1.14] shows that there is a finite field extension $L/K$, a finite set of finite places $S$ of $L$, and a model $\mathcal{X}$ for $X_{L}$ over $\mathcal{O}_{L,S}$ such that $\mathcal{X}(\mathcal{O}_{L,S})$ is not strongly thin in $X$ (i.e., $X_{L}$ has WIHP, cf. Section 1.5). We note that this (non-projective) special surface is an example of a “log-K3 surface”. In a similar vein, if $D$ is a curve of degree at most three in $\mathbb{P}^{2}_{K}$, then [Coc19, Theorem 5.2] shows that WIHP holds for the (special) surface $\mathbb{P}^{2}_{K}\setminus D$. More results on such surfaces are proven in [Coc24, Theorem 1.6], and extensions to “Campana-points” (or “semi-integral points”) on orbifold special surfaces are obtained in [NS24].

### v2

@@ Acknowledgements
The authors are grateful to Finn Bartsch, Lior Bary-Soroker, Pietro Corvaja, Pierre Dèbes, Julian Demeio, Davide Lombardo, Dan Loughran, Cedric Luger, Sebastian Petersen, Sam Streeter, and Umberto Zannier for the exchange of many ideas relevant to this survey. They are indebted to Jean-Louis Colliot-Thélène, Sebastian Petersen, Sam Streeter, Olivier Wittenberg and Umberto Zannier for numerous helpful remarks and corrections on a preliminary version.
@@ 1.1. Hilbert’s theorem
In geometric terms, this argument of Hilbert exploits that the quotient variety $\mathbb{A}^{n}_{\mathbb{Q}}/S_{n}$ is again a rational variety (i.e., birationally equivalent to $\mathbb{A}^{n}_{\mathbb{Q}}$), in fact $\mathbb{A}^{n}_{\mathbb{Q}}/S_{n}\cong\mathbb{A}^{n}_{\mathbb{Q}}$, so that the Galois cover $\mathbb{A}^{n}_{\mathbb{Q}}\rightarrow\mathbb{A}^{n}_{\mathbb{Q}}/S_{n}$ can be specialized to a Galois extension of $\mathbb{Q}$ with the same Galois group. (We will return to this geometric interpretation in Section 1.3.) In what became known as the Noether program, Emmy Noether [Noe18] started the investigation of whether $\mathbb{A}^{n}_{\mathbb{Q}}/G$ is rational for all finite groups $G\leq S_{n}$, which if true would give a strategy to solve the inverse Galois problem: Every finite group $G$ would be the Galois group of a Galois extension of $\mathbb{Q}$. However, counterexamples found by Swan [Swa69] and Voskresenskiĭ [Vos70] show that $\mathbb{A}^{n}_{\mathbb{Q}}/G$ is non-rational for $G=\mathbb{Z}/47\mathbb{Z}$, and later the same was shown for $G=\mathbb{Z}/8\mathbb{Z}$, see the references in [Len74] and [Wit23, §1.6.1] for more on the history of this problem.
@@ 1.2. Hilbertian fields
Extensive sources on Hilbertian fields are [Lan83, Chapter 9], [FJ08, Chapters 12-16], [MM18, Chapter IV], and [Völ96, Chapter 1]. For $K$ of positive characteristic, there are also variants of Definition 1.8 in the literature that are not equivalent to it. Our definition (more precisely Proposition 1.11(2)) is the one used in [FJ08] from the second edition onwards and is the one commonly used nowadays. There are also many actual strengthenings and weakenings in the literature, and we mention only some of them: fully Hilbertian fields [BP13], $g$-Hilbertian fields [FJ08, Chapter 13.6], RG-Hilbertian fields and R-Hilbertian fields [DH99], real Hilbertian fields [FHV94], $\Theta$-Hilbertian fields [FH20], and Brauer-Hilbertian fields [FSS92]. There are also the notions of Hilbertian rings [FJ08, Chapter 13.4] and integrally Hilbertian rings [BD25].
@@ 1.3. The Hilbert property
Motivated by the failure of the Noether program, Colliot-Thélène and Sansuc [CS87] investigated whether it might be possible to specialize the cover $\mathbb{A}_{\mathbb{Q}}^{n}\rightarrow\mathbb{A}_{\mathbb{Q}}^{n}/G$ to a Galois extension of $\mathbb{Q}$ with Galois group $G$ even if $\mathbb{A}^{n}_{\mathbb{Q}}/G$ is not rational. This led to the following definition by Serre [Ser08] (see Remark 1.27 below for historic details). Here, $X$ is a $K$-variety (i.e. an integral separated scheme of finite type over $K$), and a cover is a finite surjective generically étale morphism of $K$-varieties.
@@
The Hilbert property and thin sets are discussed in the textbooks [Ser08, Chapter 3], [Ser97, Chapter 9], and [FJ08, Chapter 13.5]. The first definition of HP in [CS87, p. 189] (there called “Hilbert type”) was phrased in terms of specializations of polynomials, [Ser08] introduces the notion of thin sets, while the definition we give here is adjusted for positive characteristic (cf. Remark 1.13). Some results on HP predate its formulation as a property, like [Man86, Theorem 46.1], whose first Russian edition appeared in 1972. Thin sets show up in many other places, for example as exceptional sets in modern formulations of Manin’s conjecture on the asymptotic growth of rational points of bounded height on Fano varieties, see the references in [BL19, LST22].
@@
Remark 1.29.
Another topic that we do not cover in this survey, but which is likely to become increasingly important, is the Hilbert property for algebraic stacks. Recent work has shown that the correct formulation of several arithmetic conjectures requires one to understand thin sets on stacks and not only on varieties; see for example [DY24, Section 9] and [LS25, Sections 2.2, 3.7, 10.3 and 10.4].
@@ 1.4. The weak Hilbert property
Question 1.35 ([CZ17, Question-Conjecture 2]).
@@
In fact, [CZ17] ask the question with $X$ only required to be normal, but then the answer is negative, as Remark 1.33 shows. As evidence that the answer to Question 1.35 might be positive, Corvaja and Zannier mention the case of curves (where it holds, see Example 1.31 and Section 3), conjectures of Vojta, and Manin’s conjecture in the case of Fano varieties. Like with Question 1.25, they suggest that the answer might be positive only after allowing a finite extension of the base field. We discuss this in more detail in Section 1.6.
@@ 1.5. Integral Hilbert properties
Let $K$ be a field of characteristic zero and $X$ a normal and geometrically integral $K$-variety. A positive answer to Question 1.35 would mean that if $K$ is a number field, $X$ is smooth and projective, and $X(K)$ is Zariski-dense in $X$, then $X$ has WHP, but Corvaja and Zannier considered the possibility (cf. Remark 1.36) that this might hold only potentially, in the following sense:
@@
So in the case where $K$ is a number field, a positive answer to this weaker version of Question 1.35 would offer a precise prediction when $X$ should have potential WHP, namely if and only if $X(L)$ is Zariski-dense in $X$ for some finite extension $L/K$. We now discuss for which $X$ the latter condition is expected to be satisfied. Results and conjectures in this subsection will only be used to motivate some questions and results later on.
@@
In [Cam04, Definition 2.1] (see also [Cam11, Definition 8.1] or [BJL24, Definition 1.2]) Campana introduced the class of special varieties and he conjectured that the arithmetic properties of special varieties are opposite to those conjectured by Lang for varieties of general type; see [Cam11, Conjecture 13.23]. Combined with the expected positive answer to the weakened version (see Remark 1.36) of Question 1.35, extended to finitely generated fields, this gives:
@@ 2.4. Product and fibration theorems
(1) was first published in [BFP14]; an unpublished manuscript of van den Dries from the 1990’s proves this using nonstandard methods; yet another proof was given in [BFP25, Corollary 3.5]; for $K$ a number field, (1) also follows from [HW16, Lemma 8.12]. (2) follows from [Lug24a, Theorem 1.4], generalizing the special case of smooth proper varieties in [CD+22, Theorem 1.9] due to the second-named author of this survey and Wittenberg. ∎
@@ 3. Curves
Let $K$ be Hilbertian and $X$ a curve over $K$. To see that the statement of Theorem 3.2 does not hold in this generality, recall that if $K$ is also PAC, all such $X$ have HP (Example 1.19), so the implication $\Rightarrow$ in (1)-(4) does not hold. The following Example 3.5 shows that also the implication $\Leftarrow$ in (2) does not hold. However, Proposition 3.1 implies that $\Leftarrow$ in (1) and (3) still hold, and a positive answer to Question 1.47 would imply that also $\Leftarrow$ in (4) holds in general.
@@ 4. Algebraic groups
Since connected algebraic groups are special (Example 1.45), Theorem 4.2(3) confirms a case of Conjecture 1.44. A positive answer to Question 1.47 would imply that (3) holds for arbitrary Hilbertian $K$, which in the case where $G$ is an abelian variety was confirmed for function fields $K$ in [Jav25, Theorem A]. However, Example 3.5 shows that already in the case of elliptic curves, (2) does not hold even for function fields $K$.
@@ 3. $(c)$
(1) is [Liu22, Theorem 1.3], and (2) is [CD+22, Theorem 1.4]. For cyclic $\Gamma$ and $K$ a number field, (1) in the case $G=(\mathbb{G}_{m})^{r}$ and (2) in the case $G=E^{r}$ for an elliptic curve $E$ was proven in [Zan10], which for $G=(\mathbb{G}_{m})^{r}$ built on results from [DZ07] over $\mathbb{Q}^{\rm ab}$, and for $G=E^{r}$ on ideas from [Zan00]. For a quantitative version in case (1) phrased in terms of random walks see [BG23].
@@ 5.2. Surfaces with $\kappa(X)=0$
Let $K$ be a number field, let $C$ be a smooth plane cubic in $\mathbb{P}^{2}_{K}$ and let $X=\mathbb{P}_{K}^{2}\setminus C$. Using conic fibrations on $X$, [Coc19, Theorem 1.14] shows that there is a finite field extension $L/K$, a finite set of finite places $S$ of $L$, and a model $\mathcal{X}$ for $X_{L}$ over $\mathcal{O}_{L,S}$ such that $\mathcal{X}(\mathcal{O}_{L,S})$ is not strongly thin in $X$ (i.e., $X_{L}$ has WIHP, cf. Section 1.5). We note that this (non-projective) special surface is an example of a “log-K3 surface”. In a similar vein, if $D$ is a curve of degree at most three in $\mathbb{P}^{2}_{K}$, then [Coc19, Theorem 5.2] shows that WIHP holds for the (special) surface $\mathbb{P}^{2}_{K}\setminus D$. More results on such surfaces are proven in [Coc24, Theorem 1.6] and [AL25, Theorem 1.4], and extensions to “Campana points” (or “semi-integral points”) on orbifold special surfaces are obtained in [NS24].